Assume that a disk has 256 tracks, with each track having 64 sectors and each sector is of size 4M. The cluster size in this system can be assumed to be 4 sectors. A file having the name MCS012.txt is of size 128 M. Assume that disk has 16 free - continuous clusters of 8sectors each. How can this file be allotted space on the disk? Also show the content of FAT after the space allocation to this file. You may make suitable assumptions.


1
Expert's answer
2020-09-06T17:41:28-0400

cluster_size(M) = sector_size * cluster_size(sector) = 4M * 4sectors = 16М

file_ size(M)/cluster_size(M) = 128M/16M = 8 clusters_for_file

clusters_one_tracks_= clusters_each_tracks/cluster_size(M) = 64/16M = 4 clusters on each track

Answer:

8 contiguous clusters on consecutive tracks will be allocated for the file

Directory:

File name:        MCS012.txt

First cluster number of the file:         2

 

FAT:

Number in FAT     Content

           2                     3

           3                     4

           4                     5

           5                     6

           6                     7

7                     8

           8                     9

           9                     EOF